A spherical water tower has a diameter of 20 feet. (use the value pi=3.143) Aproximately how many cubic fee of water does it contain when it is 3/4 full?
1
Expert's answer
2012-07-19T09:06:27-0400
The volume of a spherical water tower is

V = 4/3·pi·(D/2)³ = 4/3·pi·(20/2)³ = 4/3·pi·1000 feet³.

So, a tower contains

3/4·V = 3/4·4/3·pi·1000 = 1000·pi ≈ 1000·3.143 = 3143 feet³

of water when it is 3/4 full.